# and extensions of the log-Sobolev and Talagrand inequalities

### 04 Mar 2024 v1

@@ Abstract
We prove non-asymptotic error bounds for particle gradient descent (PGD) [23], a recently introduced algorithm for maximum likelihood estimation of large latent variable models obtained by discretizing a gradient flow of the free energy. We begin by showing that, for models satisfying a condition generalizing both the log-Sobolev and the Polyak–Łojasiewicz inequalities (LSI and PŁI, respectively), the flow converges exponentially fast to the set of minimizers of the free energy. We achieve this by extending a result well-known in the optimal transport literature (that the LSI implies the Talagrand inequality) and its counterpart in the optimization literature (that the PŁI implies the so-called quadratic growth condition), and applying it to our new setting. We also generalize the Bakry–Émery Theorem and show that the LSI/PŁI generalization holds for models with strongly concave log-likelihoods. For such models, we further control PGD’s discretization error, obtaining non-asymptotic error bounds. While we are motivated by the study of PGD, we believe that the inequalities and results we extend may be of independent interest.
Keywords: latent variable models, maximum marginal likelihood, gradient flows, log-Sobolev inequality, Polyak–Łojasiewicz inequality, Talagrand inequality, quadratic growth condition.
@@ 1 Introduction
if and only if $\theta_{*}$ maximizes the marginal likelihood and $q_{*}=p_{\theta_{*}}(\cdot\mid y)$; where $\operatorname{d\!}x$ in (2) denotes the Lebesgue measure and $\mathcal{M}$ the product of the parameter space $\mathbb{R}^{d_{\theta}}$ and the space $\mathcal{P}(\mathbb{R}^{d_{x}})$ probability distributions over the latent space $\mathbb{R}^{d_{x}}$ with its Euclidean Borel $\sigma$-algebra. For instance, the well-known expectation-maximization (EM) algorithm [13] can be viewed as minimizing $F$ using coordinate descent (CD) [27]. Similarly, many methods in variational inference (approximately) minimize $F$ by (i) restricting $\mathcal{P}(\mathbb{R}^{d_{x}})$ to a parametric family of distributions $(q_{\phi})_{\phi\in\Phi}$ such that the surrogate objective $(\theta,\phi)\mapsto F(\theta,q_{\phi})$ is tractable, (ii) computing a minimizer $(\theta_{*},\phi_{*})$ thereof using an appropriate optimization algorithm, and (iii) employing $(\theta_{*},q_{\phi_{*}})$ as a proxy for a genuine minimizer of $F$; e.g., see [22].
Inspired by the interpretation of the EM algorithm as CD applied to $F$, [23] examines whether it is possible to minimize $F$ over $\mathcal{M}$, rather than any restriction thereof, using analogues of optimization algorithms other than CD. In particular, they identify analogues of gradient descent (GD) applicable to $F$. To do so, they recall that the application of GD to minimizing a function $f\mathrel{\mathop{:}}\mathbb{R}^{d_{\theta}}\to\mathbb{R}$ may be viewed as the forward Euler discretization of the following ordinary differential equation (ODE) known as the gradient flow:
@@
PGD performs well in experiments [23] and is well-suited to modern computing environments. Its updates (7,8) require only evaluating $\nabla\ell\mathrel{\mathop{:}}=(\nabla_{\theta}\ell,\nabla_{x}\ell)$ and adding noise, thus facilitating its implementation in modern autodiff frameworks such as TensorFlow, PyTorch, and JAX. Moreover, each particle’s update (8) is independent of all other particles, offering scope for simple parallelisation and distribution of these steps. A full run of the algorithm can be naively implemented at a computational cost scaling as $\mathcal{O}(NK[$evaluation cost of $\nabla\ell])$, where $N$ denotes the number of particles employed and $K$ the total number of steps taken; the dependence of running time on $N$ can be further attenuated in practice through the aforementioned parallelism strategies.
Here, we validate theoretically PGD and show that, for models whose log-likelihoods $\ell$ are $\lambda$-strongly concave with $L$-Lipschitz gradients, its output, $(\Theta_{K}^{N,h},Q^{N,h}_{K})$, satisfies
@@
where $\mathsf{d}_{E}$ and $\mathsf{d}_{W_{2}}$ denote the Euclidean and Wasserstein-2 [38, Chapter 6] metrics, resp.
We begin by studying PGD’s continuous-time infinite-particle limit: the gradient flow (5). First, we show that the free energy $(F(\theta_{t},q_{t}))_{t\geq 0}$ evaluated along the flow converges at an exponential rate to its infimum,
@@
provided that the model satisfies a condition which generalizes the log-Sobolev inequality (LSI) popular in optimal transport (see e.g. [38, Chapter 21]) and the Polyak–Łojasiewicz inequality (PŁI) used in optimization (see e.g. [21]). In particular, we assume that there exists a constant $\lambda>0$ such that
@@ 1.1 Related literature
The analysis provided in [23] was limited: as explained in the ‘Our setting, notation, assumptions, rigour, and lack thereof’ paragraph of the paper’s introduction, it focused on intuition and practical application rather than theoretical validation. While the authors obtained no error bounds; they did argue with their Theorem 3 that the gradient flow (5) converges exponentially fast to $F$’s minimizer whenever $\ell$ is strongly concave. However, they imposed an unnecessary technical condition that substantially restricts the applicability of their result (that the gradient of $\ell$ be bounded uniformly).
@@
Over the last year’s course, four other works have proposed particle-based algorithms for maximum marginal likelihood estimation: [1, 24, 9, 35]. [1] consider an alternative approximation to the flow (5) which they term the ‘Interacting Particle Langevin Algorithm’ or IPLA. IPLA differs from PGD by the inclusion of an extra additive Gaussian noise term $\sqrt{2h/N}W_{k}$ in the parameter’s update (7). This difference enables an analysis of the algorithm using tools similar to those used to study the unadjusted Langevin algorithm in [17, 16]. In particular, the authors show that the $\mathcal{L}^{2}$ error of the parameter estimates produced by IPLA satisfies (9, RHS) for models satisfying the same Lipschitz gradient and strong concavity assumptions considered in the present work, which we find pleasantly consistent with some relevant results in Section 3. [24, 9] consider variants of PGD which incorporate momentum in the updates, which can outperform PGD in practice. These variants can no longer be viewed as approximations to the gradient flow (5); and consequently, require a different type of analysis. Some was conducted in [24], which established exponentially fast convergence of their algorithm’s limiting (non-gradient) flow under (11). [35] introduced two new particle-based methods for minimizing $F$. The first, ‘Stein Variational Gradient Descent EM’ or SVGD EM, is also constructed as an approximation to a gradient flow of $F$, but with respect to a different underlying geometry—the so-called ‘Stein Geometry’, as opposed to the Euclidean/Wasserstein-2 geometry underpinning (4) (whose induced distance metric is given by (10) omitting the expectations; cf. [23, Appendix A] for more on this). The second, ‘Coin EM’, involves an entirely different approach for minimizing $F$ that builds on coin betting techniques from convex optimization and obviates the need for tuning discretization step sizes. Only SVGD-EM provides a theoretical analysis, which was conducted under a natural ‘Stein’ analogue of (11): Assumption 7 in [35, Appendix A]. Let us also mention the recent work of [18], who considers, within an similar Empirical Bayes framework, a non-parametric maximum likelihood estimation problem in which they treat the prior as an unknown measure, arriving at a gradient flow that is somehow related in spirit to (5) but qualitatively different. Our results in Section 2 relate to earlier results well-known in the optimal transport and optimization literatures. We discuss these connections as we go along.
@@ 2 Convergence of the gradient flow
(i) For all $x$ in $\mathbb{R}^{d_{x}}$, $\theta\mapsto\pi_{\theta}(x)$ is differentiable; and $\theta\mapsto Z_{\theta}$ is differentiable; (ii) for all $\theta$ in $\mathbb{R}^{d_{\theta}}$, $\pi_{\theta}$ belongs to $\mathcal{P}_{2}^{1}(\mathbb{R}^{d_{x}})$; (iii) for all $\theta$ in $\mathbb{R}^{d_{\theta}}$ and $x$ in $\mathbb{R}^{d_{x}}$, $\rho_{\theta}(x)>0$;
We also assume that the gradient flow (5) has solutions in $\mathcal{M}_{2}$:
Assumption 2 (Existence of solutions).
For any initial conditions $(\theta,q)$ in $\mathcal{M}_{2}$, (5) has a weak solution $(\theta_{t},q_{t})_{t\geq 0}$ with $(\theta_{0},q_{0})=(\theta,q)$. By a weak solution, we mean a function mapping from $[0,\infty)$ to $\mathcal{M}_{2}$ such that for all $t\geq 0$ and twice continuously differentiable and compactly supported functions $\phi\mathrel{\mathop{:}}\mathbb{R}^{d_{x}}\to\mathbb{R}$,
$\frac{\operatorname{d\!}\theta_{t}}{\operatorname{d\!}t}$ $=\int\nabla_{\theta}\ell(\theta_{t},x)q_{t}(\operatorname{d\!}x),$
$\frac{\operatorname{d\!}}{\operatorname{d\!}t}\int\phi(x)q_{t}(\operatorname{d\!}x)$ $=\int\Delta_{x}\phi(x)q_{t}(\operatorname{d\!}x)-\int\left\langle{\nabla_{x}\ell(\theta_{t},x)},{\nabla_{x}\phi(x)}\right\rangle q_{t}(\operatorname{d\!}x);$
[…]
We verify in Section 3 that the above holds whenever the model’s log-likelihood has a Lipschitz gradient. Lastly, we assume that the solutions are sufficiently regular:
@@
Let $(\theta_{t},q_{t})_{t\geq 0}$ be any weak solution of (5). Then, for all $t>0$, $(\theta_{t},q_{t})\in\mathcal{M}_{2}^{1}$ and
$\frac{\operatorname{d\!}F}{\operatorname{d\!}t}(\theta_{t},q_{t})=-I(\theta_{t},q_{t})\quad\forall t>0.$ (15)
Formally arguing (15) is straightforward (cf. Appendix A.1). In fact, we believe that the above assumption holds under regularity conditions on the model of the sort in Assumption 1. However, rigorously arguing this is technically involved and lies beyond the scope of the paper. In the special case that $\pi_{\theta}$ is a single distribution $\pi$ independent of $\theta$, the flow (5) essentially reduces to the following well-known Fokker-Planck equation,
@@
whose solutions’ regularity properties are well-understood (e.g., [20, Theorem 5.1] shows that the solutions are smooth whenever $\log(\pi)$ is smooth). In this case, (15) reduces to de Bruijn’s identity which is known to hold whenever $\pi$ is sufficiently regular; e.g. see [5, Proposition 5.2.2] or [31, Lemma 1].
@@ 2.1 The extended log-Sobolev inequality and $(F(\theta_{t},q_{t}))_{t\geq 0}$’s convergence
If Assumptions 13 hold, $(\theta_{0},q_{0})$ belongs to $\mathcal{M}_{2}$, and the measures $(\rho_{\theta}(\operatorname{d\!}x))_{\theta\in\mathbb{R}^{d_{\theta}}}$ satisfy the xLSI with constant $\lambda$ (Definition 1), then
@@
The leftmost inequality in (17) follows from the definition of $F_{*}$. The rightmost equality follows by combining the xLSI (11) with de Bruijn’s identity (15) and applying Grönwall’s inequality. ∎
@@
We refer to the inequality in Definition 1 as the ‘xLSI’ because it extends the classical log-Sobolev inequality (LSI); e.g., see [38, Definition 21.1]. While the xLSI is a statement about a parametrized family of measures, the LSI is a statement about a single probability measure. In particular, if $\pi_{\theta}$ is a distribution $\pi$ independent of $\theta$, then (11) reads
@@
For this reason, Theorem 1 extends a well-known result (e.g., [5, Section 5.2]) stating that, if $\pi$ satisfies the LSI (20) and $(q_{t})_{t\geq 0}$ solves the Fokker-Planck equation (16), then the $\mathrm{KL}$ divergence from $q_{t}$ to $\pi$ decays exponentially fast.
@@
setting $q\mathrel{\mathop{:}}=\pi_{\theta}$ in (11), we recover (21) with $f(\theta)\mathrel{\mathop{:}}=-\log(Z_{\theta})$. (The exchange of limits in (22) is valid whenever Fisher’s identity holds, which is a common assumption in the study of latent variable models and the EM algorithm; it is satisfied whenever e.g. $\pi_{\theta}$ belongs to an exponential family [15, Appendix D].) In this case, Theorem 1 reduces to a well-known result showing that $f$ converges exponentially fast along the Euclidean gradient flow (3) to its infimum whenever $f$ satisfies the PŁI; e.g., see [37, Proposition 2.3].
Among other reasons, the PŁI (21) is popular in optimization because—without requiring the objective to be convex—it implies that the objective’s stationary points are global minimizers. In our case, the xLSI implies that the marginal likelihood’s stationary points are global maximizers.

### 11 Apr 2024 v2

@@ Abstract
We prove non-asymptotic error bounds for particle gradient descent (PGD) [26], a recently introduced algorithm for maximum likelihood estimation of large latent variable models obtained by discretizing a gradient flow of the free energy. We begin by showing that, for models satisfying a condition generalizing both the log-Sobolev and the Polyak–Łojasiewicz inequalities (LSI and PŁI, respectively), the flow converges exponentially fast to the set of minimizers of the free energy. We achieve this by extending a result well-known in the optimal transport literature (that the LSI implies the Talagrand inequality) and its counterpart in the optimization literature (that the PŁI implies the so-called quadratic growth condition), and applying it to our new setting. We also generalize the Bakry–Émery Theorem and show that the LSI/PŁI generalization holds for models with strongly concave log-likelihoods. For such models, we further control PGD’s discretization error, obtaining non-asymptotic error bounds. While we are motivated by the study of PGD, we believe that the inequalities and results we extend may be of independent interest. Keywords: latent variable models, maximum marginal likelihood, gradient flows, log-Sobolev inequality, Polyak–Łojasiewicz inequality, Talagrand inequality, quadratic growth condition.
@@ 1 Introduction
if and only if $\theta_{*}$ maximizes the marginal likelihood and $q_{*}=p_{\theta_{*}}(\cdot\mid y)$; where $\operatorname{d\!}x$ in (2) denotes the Lebesgue measure and $\mathcal{M}$ the product of the parameter space $\mathbb{R}^{d_{\theta}}$ and the space $\mathcal{P}(\mathbb{R}^{d_{x}})$ probability distributions over the latent space $\mathbb{R}^{d_{x}}$ with its Euclidean Borel $\sigma$-algebra. For instance, the well-known expectation-maximization (EM) algorithm [15] can be viewed as minimizing $F$ using coordinate descent (CD) [32]. Similarly, many methods in variational inference (approximately) minimize $F$ by (i) restricting $\mathcal{P}(\mathbb{R}^{d_{x}})$ to a parametric family of distributions $(q_{\phi})_{\phi\in\Phi}$ such that the surrogate objective $(\theta,\phi)\mapsto F(\theta,q_{\phi})$ is tractable, (ii) computing a minimizer $(\theta_{*},\phi_{*})$ thereof using an appropriate optimization algorithm, and (iii) employing $(\theta_{*},q_{\phi_{*}})$ as a proxy for a genuine minimizer of $F$; e.g., see [24]. Inspired by the interpretation of the EM algorithm as CD applied to $F$, [26] examines whether it is possible to minimize $F$ over $\mathcal{M}$, rather than any restriction thereof, using analogues of optimization algorithms other than CD. In particular, they identify analogues of gradient descent (GD) applicable to $F$. To do so, they recall that the application of GD to minimizing a function $f\mathrel{\mathop{:}}\mathbb{R}^{d_{\theta}}\to\mathbb{R}$ may be viewed as the forward Euler discretization of the following ordinary differential equation (ODE) known as the gradient flow:
@@
PGD performs well in experiments [26] and is well-suited to modern computing environments. Its updates (7,8) require only evaluating $\nabla\ell\mathrel{\mathop{:}}=(\nabla_{\theta}\ell,\nabla_{x}\ell)$ and adding noise, thus facilitating its implementation in modern autodiff frameworks such as TensorFlow, PyTorch, and JAX. Moreover, each particle’s update (8) is independent of all other particles, offering scope for simple parallelisation and distribution of these steps. A full run of the algorithm can be naively implemented at a computational cost scaling as $\mathcal{O}(NK[$evaluation cost of $\nabla\ell])$, where $N$ denotes the number of particles employed and $K$ the total number of steps taken; the dependence of running time on $N$ can be further attenuated in practice through the aforementioned parallelism strategies. Here, we validate theoretically PGD and show that, for models whose log-likelihoods $\ell$ are $\lambda$-strongly concave with $L$-Lipschitz gradients, its output, $(\Theta_{K}^{N,h},Q^{N,h}_{K})$, satisfies
@@
where $\mathsf{d}_{E}$ and $\mathsf{d}_{W_{2}}$ denote the Euclidean and Wasserstein-2 [45, Chapter 6] metrics, resp. We begin by studying PGD’s continuous-time infinite-particle limit: the gradient flow (5). First, we show that the free energy $(F(\theta_{t},q_{t}))_{t\geq 0}$ evaluated along the flow converges at an exponential rate to its infimum,
@@
provided that the model satisfies a condition which generalizes the log-Sobolev inequality (LSI) popular in optimal transport (see e.g. [45, Chapter 21]) and the Polyak–Łojasiewicz inequality (PŁI) used in optimization (see e.g. [23]). In particular, we assume that there exists a constant $\lambda>0$ such that
@@ 1.1 Related literature
The analysis provided in [26] was limited: as explained in the ‘Our setting, notation, assumptions, rigour, and lack thereof’ paragraph of the paper’s introduction, it focused on intuition and practical application rather than theoretical validation. While the authors obtained no error bounds; they did argue with their Theorem 3 that the gradient flow (5) converges exponentially fast to $F$’s minimizer whenever $\ell$ is strongly concave. However, they imposed an unnecessary technical condition that substantially restricts the applicability of their result (that the gradient of $\ell$ be bounded uniformly). Over the last year’s course, five other works have proposed particle-based algorithms for maximum marginal likelihood estimation: [2, 28, 11, 40, 20]. [2] consider an alternative approximation to the flow (5) which they term the ‘Interacting Particle Langevin Algorithm’ or IPLA. IPLA differs from PGD by the inclusion of an extra additive Gaussian noise term $\sqrt{2h/N}W_{k}$ in the parameter’s update (7). This difference enables an analysis of the algorithm using tools similar to those used to study the unadjusted Langevin algorithm in [19, 18]. In particular, the authors show that the $\mathcal{L}^{2}$ error of the parameter estimates produced by IPLA satisfies (9, RHS) for models satisfying the same Lipschitz gradient and strong concavity assumptions considered in the present work, which we find pleasantly consistent with some relevant results in Section 3. [28, 11] consider variants of PGD which incorporate momentum in the updates, which can outperform PGD in practice. These variants can no longer be viewed as approximations to the gradient flow (5); and consequently, require a different type of analysis. Some was conducted in [28], which established exponentially fast convergence of their algorithm’s limiting (non-gradient) flow under (11). [40] introduced two new particle-based methods for minimizing $F$. The first, ‘Stein Variational Gradient Descent EM’ or SVGD EM, is also constructed as an approximation to a gradient flow of $F$, but with respect to a different underlying geometry—the so-called ‘Stein Geometry’, as opposed to the Euclidean/Wasserstein-2 geometry underpinning (4) (whose induced distance metric is given by (10) omitting the expectations; cf. [26, Appendix A] for more on this). The second, ‘Coin EM’, involves an entirely different approach for minimizing $F$ that builds on coin betting techniques from convex optimization and obviates the need for tuning discretization step sizes. Only SVGD-EM provides a theoretical analysis, which was conducted under a natural ‘Stein’ analogue of (11): Assumption 7 in [40, Appendix A]. [20] considers a non-parametric maximum likelihood estimation problem in which they treat the prior as an unknown measure, arriving at a gradient flow that is related in spirit to (5) but qualitatively different.
Our results in Section 2 relate to earlier results well-known in the optimal transport and optimization literatures. We discuss these connections as we go along.
@@ 2 Convergence of the gradient flow
(i) For all $x$ in $\mathbb{R}^{d_{x}}$, $\theta\mapsto\pi_{\theta}(x)$ is differentiable; and $\theta\mapsto Z_{\theta}$ is differentiable; (ii) for all $\theta$ in $\mathbb{R}^{d_{\theta}}$, $\pi_{\theta}$ is twice continuously differentiable; (iii) for all $\theta$ in $\mathbb{R}^{d_{\theta}}$ and $x$ in $\mathbb{R}^{d_{x}}$, $\rho_{\theta}(x)>0$;
Let $\mathcal{C}^{i}$ denote the space of $i$-times continuously differentiable functions from some space to another that will be specified in the brackets that follow. $\mathcal{C}^{i}_{c}$ indicates that such functions are taken to have compact support. $\mathcal{C}^{i,j}$ denote the functions that $\mathcal{C}^{i}$ in the first variable and $\mathcal{C}^{j}$ in the second. The next assumptions is on the existence of suitable solutions $(\theta_{t},q_{t})_{t\geq 0}$ to the gradient flow (5).
@@
For any initial conditions $(\theta,q)$ in $\mathcal{M}_{2}$, (5) has a classical solution $(\theta_{t},q_{t})_{t\geq 0}$ with $(\theta_{0},q_{0})=(\theta,q)$. Furthermore, for all $t\geq 0$, $q_{t}$ has a Lebesgue density in $\mathcal{C}^{1,2}([0,\infty)\times\mathbb{R}^{d_{x}},\mathbb{R}^{+})$ and $\theta_{t}\in\mathcal{C}^{1}([0,\infty),\mathbb{R}^{d_{\theta}})$.
We verify in Section 3 that the above holds whenever the model’s log-likelihood $\ell$ has a Lipschitz gradient, but we believe softer smoothness assumptions on $\ell$ should suffice. In the special case that $\pi_{\theta}$ is a single distribution $\pi$ independent of $\theta$, the flow (5) essentially reduces to the following well-known Fokker-Planck equation,
@@
whose solutions’ regularity properties are well-understood (e.g., [22, Theorem 5.1] shows that the solutions are smooth whenever $\log(\pi)$ is smooth).
@@ 2.1 The extended log-Sobolev inequality and $(F(\theta_{t},q_{t}))_{t\geq 0}$’s convergence
If Assumptions 12 hold, $(\theta_{0},q_{0})$ belongs to $\mathcal{M}_{2}$, and the measures $(\rho_{\theta}(\operatorname{d\!}x))_{\theta\in\mathbb{R}^{d_{\theta}}}$ satisfy the xLSI with constant $\lambda$, then
@@
The leftmost inequality in (16) follows from the definition of $F_{*}$. As we show in Lemma 13 in Appendix A.1, under our assumptions there holds the following extension to de Bruijin’s identity:
$\frac{\operatorname{d\!}}{\operatorname{d\!}t}F(\theta_{t},q_{t})=-I(\theta_{t},q_{t})\quad\forall t>0,$ (17)
(cf. [6, Proposition 5.2.2]). Now the rightmost equality follows by combining the xLSI (11) with (17) and applying Grönwall’s inequality. ∎
@@
We refer to the inequality in Definition 1 as the ‘xLSI’ because it extends the classical log-Sobolev inequality (LSI); e.g., see [45, Definition 21.1]. While the xLSI is a statement about a parametrized family of measures, the LSI is a statement about a single probability measure. In particular, if $\pi_{\theta}$ is a distribution $\pi$ independent of $\theta$, then (11) reads
@@
For this reason, Theorem 1 extends a well-known result (e.g., [6, Section 5.2]) stating that, if $\pi$ satisfies the LSI (20) and $(q_{t})_{t\geq 0}$ solves the Fokker-Planck equation (15), then the $\mathrm{KL}$ divergence from $q_{t}$ to $\pi$ decays exponentially fast.
@@
setting $q\mathrel{\mathop{:}}=\pi_{\theta}$ in (11), we recover (21) with $f(\theta)\mathrel{\mathop{:}}=-\log(Z_{\theta})$. (The exchange of limits in (22) is valid whenever Fisher’s identity holds, which is a common assumption in the study of latent variable models and the EM algorithm; it is satisfied whenever e.g. $\pi_{\theta}$ belongs to an exponential family [17, Appendix D].) In this case, Theorem 1 reduces to a well-known result showing that $f$ converges exponentially fast along the Euclidean gradient flow (3) to its infimum whenever $f$ satisfies the PŁI; e.g., see [44, Proposition 2.3]. Among other reasons, the PŁI (21) is popular in optimization because—without requiring the objective to be convex—it implies that the objective’s stationary points are global minimizers. In our case, the xLSI implies that the marginal likelihood’s stationary points are global maximizers.